# K theoretic virtual fundamental cycle of global Kuranishi chart

### v1

@@ 3.1. The Cap product
In this Subsection, we define the K-theoretic cap product on $\mathcal{Y}$. Given $(H,\phi,T)\in K^{fake}_{0}(\mathcal{Y})$ and $(\mathcal{E},\alpha)\in K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})$, we construct an element $(\tilde{H},\tilde{\phi},\tilde{T})\in K^{fake}_{0}(\mathcal{Y})$ as the cap product $\alpha\cap T$ following [8].
We first describe the Hilbert $\mathbb{C}(\mathcal{Y})$-module structure $(\tilde{H},\tilde{\phi})$. For a complex orbi-bundle $\mathcal{E}$, let $C(\mathcal{E})$ be continuous sections of $\mathcal{E}$. Then $C(\mathcal{E})$ is a finitely generated projective module over $\mathbb{C}(\mathcal{Y})$. So there is an idempotent matrices $A$ of size $n$, with entries in $\mathbb{C}(\mathcal{Y})$, whose range is $C(\mathcal{E})$. It is viewed as a map $A:H^{n}\to H^{n}$ via $\phi:\mathbb{C}(\mathcal{Y})\to L(H)$. Set $\tilde{H}=\operatorname{im}A$ and $\tilde{\phi}$ be the natural $\mathbb{C}(\mathcal{Y})$-module structures on $\tilde{H}$.
To describe the Fredholm operator $\tilde{T}:\tilde{H}\to\tilde{H}$, let $C(\alpha)$ be the $\mathbb{C}(\mathcal{Y})$-module homomorphism $C(\mathcal{E})\to C(\mathcal{E})$ induced by $\alpha:\mathcal{E}\to\mathcal{E}$. Let $T\otimes^{\prime}1,1\otimes C(\alpha)$ be the induced maps on $\operatorname{im}A$, and
@@
    We have $\tilde{H}=H\otimes_{\mathbb{C}(\mathcal{Y})}C(\mathcal{E})$ as vector spaces, but different $A$’s define on this vector space a different inner product. Nonetheless, all such inner products define the same topology.
@@
    While $1\otimes C(\alpha)$ is the actual tensor product, we emphasize that the ‘tensor product’ $T\otimes 1$ does not make sense, since the (pseudo-differential) operator $T$ is not $\mathbb{C}(\mathcal{Y})$-linear. Therefore, we use the notation $\otimes^{\prime}$ and need to introduce the projection $A$ in our construction.
@@
$\begin{array}[]{ll}[\tilde{T},\tilde{\phi}(f)]&=[T,\phi(f)]\otimes^{\prime}1;\\ \tilde{T}^{2}&=T^{2}\otimes^{\prime}1+1\otimes C(\alpha)^{2}.\end{array}$
Hence $\tilde{T}$ almost commutes with $\tilde{\phi}(f)$ and $\tilde{T}^{\ast}=\tilde{T}$ has finite dimensional kernel. ∎
@@
$(H,\phi,T)\in K^{fake}_{0}(\mathcal{Y})$ only depends on $T,\phi$ and the topology on $H$ (and is independent of the inner products on $H$).
@@
Let $(H,\phi,T)$ be a Kasparov module. Let $H^{\prime}=H$ as $\mathbb{C}(\mathcal{Y})$-modules and topological spaces, each with a certain inner product. Then
$\tilde{T}_{t}=\left[\begin{matrix}T\cos(\frac{\pi}{2}t)&\sin(\frac{\pi}{2}t)I\\ \sin(\frac{\pi}{2}t)I&-T^{\prime}\cos(\frac{\pi}{2}t)\end{matrix}\right]$
defines a homotopy from $(H,\phi,T)\ominus(H^{\prime},\phi,T)$ to a degenerate Kasparov module. So $(H^{\prime},\phi,T)=(H,\phi,T)$ in $K^{fake}_{0}(\mathcal{Y})$. ∎
@@
The (fake) cap product is a map $K_{0}^{fake}(\mathcal{Y})\otimes K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})\to K_{0}^{fake}(\mathcal{Y})$.
@@
    The cap product $(\tilde{H},\tilde{\phi},\tilde{T})$ is independent of $A$ (and $n$) chosen.
@@
Let $A,A^{\prime}$ be idempotent matrices of sizes $n,n^{\prime}$ with $\mathbb{C}(\mathcal{Y})$ coefficients, whose images are identified via a $\mathbb{C}(\mathcal{Y})$-module homomorphism $\iota$ that is also an isometry. Then the composition
@@
is a lift of $\iota$ as a module homomorphism, thus realizing the map $\tilde{H}\to\tilde{H}^{\prime}$ as a matrix transformation, and so this map is continuous. Likewise the map $\tilde{H}^{\prime}\to\tilde{H}$ induced by $\iota^{-1}$ is also continuous, so the topologies on $\tilde{H}$ and $\tilde{H}^{\prime}$ are identical. Moreover, $\tilde{\iota}$ intertwines $\phi(f)$’s and $T$’s up to a compact operator, so different $A$’s give unitarily equivalent cap products.
Different representatives of $K$-homology differ by a sequence of unitary equivalence, homotopy, or direct sum with degenerate objects. Different representatives of the same element in the $K$ group differ by a sequence of isomorphism, homotopy, or direct sum with degenerated objects. The cap product, in turn, is also connected by a sequence of unitary equivalence, homotopy, or direct sum with degenerated objects. So the cap product
@@
Given an inclusion of compact sets $\mathcal{Y}\subseteq\mathcal{Y}^{\prime}$, we viewed $C(\mathcal{E}|_{\mathcal{Y}})$ as the range of a $\mathbb{C}(\mathcal{Y})$-matrix. Indeed, restricting the coefficients of this matrix (that are functions on $\mathcal{Y}^{\prime}$) to $\mathcal{Y}$, we obtain a matrix whose range gives $C(\mathcal{E}|_{\mathcal{Y}})$. Computing under this specific choice of basis shows that each inclusion of $\mathcal{Y}$ induces a map on cap products. ∎
@@ 3.2. $K^{0}(\mathcal{Y},\partial\mathcal{Y})$-module structures
In this Subsection, we show that the $K^{0}(\mathcal{Y},\partial\mathcal{Y})$ module structures on $K^{0}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})$ and $K_{0}(\mathcal{Y})$ are intertwined by $\mu_{\mathcal{Y}}$.
Let $\pi:B^{\ast}\mathcal{Y}\to\mathcal{Y}$. Given elements $(\mathcal{E},\alpha)\in K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})$ and $(\mathcal{F},\beta)\in K^{0,fake}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})$, we define ([4] page 490)
@@
This element is supported in the (compact) intersection of $\operatorname{supp}\alpha\subseteq\mathcal{Y}\subseteq B^{\ast}\mathcal{Y}$ and $\operatorname{supp}\beta\subseteq B^{\ast}\mathcal{Y}$, making $K^{0,fake}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})$ a $K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})$ module. On the other hand, $K^{fake}_{0}(\mathcal{Y})$ carries a $K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})$-module structure given by the cap product. We use $\hat{\otimes}$ for the completion of the algebraic tensor product.
@@
Let $\mathcal{E},\mathcal{F}$ be Hermitian bundles with trivial isotropy group action over a compact orbifold $\mathcal{Y}$. Then $L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\mathcal{F})\cong L^{2}(\mathcal{E}\otimes_{\mathbb{C}}\mathcal{F})$ as $\mathbb{C}(\mathcal{Y})$-modules and as topological spaces.
@@
If $\mathcal{F}=\underline{n}$, the $n$-dimensional trivial bundle over $\mathcal{Y}$, $C(\underline{n})=(\mathbb{C}(\mathcal{Y}))^{n}$ induces an isomorphism
$L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\underline{n})=(L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}\mathbb{C}(\mathcal{Y}))^{n}=(L^{2}(\mathcal{E}\otimes\underline{1}))^{n}=L^{2}(\mathcal{E}\otimes\underline{n}),$
as $\mathbb{C}(\mathcal{Y})$-modules and topological spaces.
For general $\mathcal{F}$, there is an $n$ and a complex vector bundle $\mathcal{F}^{\prime}$ such that $\mathcal{F}\oplus\mathcal{F}^{\prime}=\underline{n}$ by the compactness of $\mathcal{Y}$. The natural map $L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\mathcal{F})\to L^{2}(\mathcal{E}\otimes\underline{n})$ is injective with image contained in $L^{2}(\mathcal{E}\otimes\mathcal{F})$, since $C(\underline{n})=C(\mathcal{F})\oplus C(\mathcal{F}^{\prime})$. So $L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\mathcal{F})$ is viewed as a sub-module of $L^{2}(\mathcal{E}\otimes\mathcal{F})$. Similarly $L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\mathcal{F}^{\prime})\subseteq L^{2}(\mathcal{E}\otimes\mathcal{F}^{\prime})$. Now, it follows from
@@
$L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}\mathbb{C}(\mathcal{F})=L^{2}(\mathcal{E}\otimes\mathcal{F}).$
@@
To do this, let $\beta:\mathcal{F}\to\mathcal{F}$ be an element in $K^{0,fake}(\mathcal{Y})$. Let $P:L^{2}(\mathcal{E})\to L^{2}(\mathcal{E})$ be a pseudo-differential operator of order 00 that represents $\alpha:\mathcal{E}\to\mathcal{E}$. Let $\mathcal{U}$ be a relatively compact open sub-manifold, such that $\alpha$ is an isomorphism, and $P$ is multiplication by a function outside $\mathcal{U}$. We have the following results.
$\begin{array}[]{lll}\beta\boxtimes\sigma(P)&=\beta\otimes 1-\varepsilon\otimes\sigma(P):&\pi^{\ast}\mathcal{E}\otimes\pi^{\ast}\mathcal{F}\to\pi^{\ast}\mathcal{E}\otimes\pi^{\ast}\mathcal{F}\\ P\cap\beta&=P\otimes 1-\varepsilon\otimes C(\beta):&L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}\mathbb{C}(\mathcal{F})\to L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}\mathbb{C}(\mathcal{F})\end{array}$
@@
represents $\beta\boxtimes\sigma(P)$ because $\sigma(P\otimes 1)=\sigma(P)\otimes 1$. It also represents $[P]\cap\beta$ by the previous lemma. Therefore $\mu([\sigma(P)]\boxtimes[\beta])=[P]\cap[\beta]$, and so the diagram commute. ∎
3.3. K-homological maps $i^{!}$ and $i_{\ast}$

### v2

@@ 3.1. The Cap product
In this Subsection, we define the K-theoretic cap product on $\mathcal{Y}$. Given $(H,\phi,T)\in K^{fake}_{0}(\mathcal{Y})$ and $(\mathcal{E},\alpha)\in K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})$, we construct an element $(\tilde{H},\tilde{\phi},\tilde{T})\in K^{fake}_{0}(\mathcal{Y})$ as the cap product $\alpha\cap T$, following [8].
We first describe the Hilbert $\mathbb{C}(\mathcal{Y})$-module structure $(\tilde{H},\tilde{\phi})$. For a complex orbi-bundle $\mathcal{E}$, let $C(\mathcal{E})$ be continuous sections of $\mathcal{E}$. Then $C(\mathcal{E})$ is a finitely generated projective module over $\mathbb{C}(\mathcal{Y})$. Thus, there is an idempotent matrix $A$ of size $n$, with entries in $\mathbb{C}(\mathcal{Y})$, whose range is $C(\mathcal{E})$. It is viewed as a map $A:H^{n}\to H^{n}$ via $\phi:\mathbb{C}(\mathcal{Y})\to L(H)$. Set $\tilde{H}=\operatorname{im}A$ and let $\tilde{\phi}$ be the natural $\mathbb{C}(\mathcal{Y})$-module structure on $\tilde{H}$.
To describe the Fredholm operator $\tilde{T}:\tilde{H}\to\tilde{H}$, let $C(\alpha)$ be the $\mathbb{C}(\mathcal{Y})$-module homomorphism $C(\mathcal{E})\to C(\mathcal{E})$ induced by $\alpha:\mathcal{E}\to\mathcal{E}$. Let $T\otimes^{\prime}1$ and $1\otimes C(\alpha)$ be the induced maps on $\operatorname{im}A$, and define
@@
Recall that in the introduction, we defined $\varepsilon$ as the operator on $\mathbb{Z}_{2}$-graded objects acting as $(-1)^{a}$ on the direct summand of degree $a$.
@@
    We have $\tilde{H}=H\otimes_{\mathbb{C}(\mathcal{Y})}C(\mathcal{E})$ as vector spaces, but different $A$’s define different inner products on this vector space. Nonetheless, all such inner products define the same topology (which is all we need by Lemma 3.2).
@@
    While $1\otimes C(\alpha)$ is the actual tensor product, we emphasize that the ‘tensor product’ $T\otimes 1$ does not make sense since the (pseudo-differential) operator $T$ is not $\mathbb{C}(\mathcal{Y})$-linear. Therefore, we use the notation $\otimes^{\prime}$ and need to introduce the projection $A$ in our construction.
@@
$\begin{array}[]{ll}[\tilde{T},\tilde{\phi}(f)]&=[T,\phi(f)]\otimes^{\prime}1;\\[2.84526pt] \tilde{T}^{2}&=T^{2}\otimes^{\prime}1+1\otimes C(\alpha)^{2}.\end{array}$
Hence, $\tilde{T}$ almost commutes with $\tilde{\phi}(f)$ and $\tilde{T}^{\ast}=\tilde{T}$ has a finite-dimensional kernel. ∎
@@
$(H,\phi,T)\in K^{fake}_{0}(\mathcal{Y})$ depends only on $T$, $\phi$, and the topology on $H$ (and is independent of the inner products on $H$).
@@
Let $(H,\phi,T)$ be a Kasparov module. Let $H^{\prime}=H$ as $\mathbb{C}(\mathcal{Y})$-modules and topological spaces, each endowed with a certain inner product. Then
$\tilde{T}_{t}=\left[\begin{matrix}T\cos(\frac{\pi}{2}t)&\sin(\frac{\pi}{2}t)I\\[2.84526pt] \sin(\frac{\pi}{2}t)I&-T\cos(\frac{\pi}{2}t)\end{matrix}\right]$
defines a homotopy from $(H,\phi,T)\ominus(H^{\prime},\phi,T)$ to a degenerate Kasparov module. Thus, $(H^{\prime},\phi,T)=(H,\phi,T)$ in $K^{fake}_{0}(\mathcal{Y})$. ∎
@@
The (fake) cap product is a map
$K_{0}^{fake}(\mathcal{Y})\otimes K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})\to K_{0}^{fake}(\mathcal{Y}).$
@@
    The cap product $(\tilde{H},\tilde{\phi},\tilde{T})$ is independent of the choice of $A$ (and $n$).
@@
Let $A,A^{\prime}$ be idempotent matrices of sizes $n$ and $n^{\prime}$ with $\mathbb{C}(\mathcal{Y})$ coefficients, whose images are identified via a $\mathbb{C}(\mathcal{Y})$-module homomorphism $\iota$ that is also an isometry. Then the composition
@@
is a lift of $\iota$ as a module homomorphism, thus realizing the map $\tilde{H}\to\tilde{H}^{\prime}$ as a matrix transformation, and so this map is continuous. Likewise, the map $\tilde{H}^{\prime}\to\tilde{H}$ induced by $\iota^{-1}$ is also continuous, so the topologies on $\tilde{H}$ and $\tilde{H}^{\prime}$ are identical. Moreover, $\tilde{\iota}$ intertwines $\phi(f)$’s and $T$’s up to a compact operator, so different $A$’s give unitarily equivalent cap products.
Different representatives of $K$-homology differ by a sequence of unitary equivalences, homotopies, or direct sums with degenerate objects. Similarly, different representatives of the same element in the $K$-group differ by a sequence of isomorphisms, homotopies, or direct sums with degenerate objects. The cap product, in turn, is also connected by a sequence of unitary equivalences, homotopies, or direct sums with degenerate objects. Thus, the cap product
@@
Given an inclusion of compact sets $\mathcal{Y}\subseteq\mathcal{Y}^{\prime}$, we view $C(\mathcal{E}|_{\mathcal{Y}})$ as the range of a $\mathbb{C}(\mathcal{Y})$-matrix. Indeed, by restricting the coefficients of this matrix (which are functions on $\mathcal{Y}^{\prime}$) to $\mathcal{Y}$, we obtain a matrix whose range gives $C(\mathcal{E}|_{\mathcal{Y}})$. Computing under this specific choice of basis shows that each inclusion of $\mathcal{Y}$ induces a map on cap products. ∎
@@ 3.2. $K^{0}(\mathcal{Y},\partial\mathcal{Y})$-module structures
In this subsection, we show that the $K^{0}(\mathcal{Y},\partial\mathcal{Y})$ module structures on $K^{0}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})$ and on $K_{0}(\mathcal{Y})$ are intertwined by $\mu_{\mathcal{Y}}$.
Let $\pi:B^{\ast}\mathcal{Y}\to\mathcal{Y}$. Given elements $(\mathcal{E},\alpha)\in K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})$ and $(\mathcal{F},\beta)\in K^{0,fake}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})$, we define (see [4] page 490)
@@
This element is supported in the (compact) intersection of $\operatorname{supp}\alpha\subseteq\mathcal{Y}\subseteq B^{\ast}\mathcal{Y}$ and $\operatorname{supp}\beta\subseteq B^{\ast}\mathcal{Y}$, making $K^{0,fake}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})$ a $K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})$-module. On the other hand, $K^{fake}_{0}(\mathcal{Y})$ carries a $K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})$-module structure given by the cap product. We use $\hat{\otimes}$ for the completion of the algebraic tensor product.
@@
Let $\mathcal{E},\mathcal{F}$ be Hermitian bundles with trivial isotropy group action over a compact orbifold $\mathcal{Y}$. Then
$L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\mathcal{F})\cong L^{2}(\mathcal{E}\otimes_{\mathbb{C}}\mathcal{F})$
as $\mathbb{C}(\mathcal{Y})$-modules and as topological spaces.
@@
If $\mathcal{F}=\underline{n}$, the $n$-dimensional trivial bundle over $\mathcal{Y}$, then $C(\underline{n})=(\mathbb{C}(\mathcal{Y}))^{n}$ induces an isomorphism $L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\underline{n})=\Bigl{(}L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}\mathbb{C}(\mathcal{Y})\Bigr{)}^{n}=\Bigl{(}L^{2}(\mathcal{E}\otimes\underline{1})\Bigr{)}^{n}=L^{2}(\mathcal{E}\otimes\underline{n})$, as $\mathbb{C}(\mathcal{Y})$-modules and topological spaces.
For general $\mathcal{F}$, there is an $n$ and a complex vector bundle $\mathcal{F}^{\prime}$ such that $\mathcal{F}\oplus\mathcal{F}^{\prime}=\underline{n}$ by the compactness of $\mathcal{Y}$. The natural map $L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\mathcal{F})\to L^{2}(\mathcal{E}\otimes\underline{n})$ is injective with image contained in $L^{2}(\mathcal{E}\otimes\mathcal{F})$, since $C(\underline{n})=C(\mathcal{F})\oplus C(\mathcal{F}^{\prime})$. Thus, $L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\mathcal{F})$ may be viewed as a sub-module of $L^{2}(\mathcal{E}\otimes\mathcal{F})$. Similarly, $L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\mathcal{F}^{\prime})\subseteq L^{2}(\mathcal{E}\otimes\mathcal{F}^{\prime})$. Now, it follows from
@@
$L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\mathcal{F})=L^{2}(\mathcal{E}\otimes\mathcal{F}).$
@@
To do this, let $\beta:\mathcal{F}\to\mathcal{F}$ be an element in $K^{0,fake}(\mathcal{Y})$. Let
$P:L^{2}(\mathcal{E})\to L^{2}(\mathcal{E})$
be a pseudo-differential operator of order 00 that represents
$\alpha:\mathcal{E}\to\mathcal{E}.$
Let $\mathcal{U}$ be a relatively compact open sub-manifold such that $\alpha$ is an isomorphism and such that $P$ is given by multiplication by a function outside $\mathcal{U}$. We have the following results.
$\begin{array}[]{lll}\beta\boxtimes\sigma(P)&=\beta\otimes 1-\varepsilon\otimes\sigma(P):&\pi^{\ast}\mathcal{E}\otimes\pi^{\ast}\mathcal{F}\to\pi^{\ast}\mathcal{E}\otimes\pi^{\ast}\mathcal{F},\\ P\cap\beta&=P\otimes 1-\varepsilon\otimes C(\beta):&L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\mathcal{F})\to L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\mathcal{F}).\end{array}$
@@
represents $\beta\boxtimes\sigma(P)$ because $\sigma(P\otimes 1)=\sigma(P)\otimes 1$. It also represents $[P]\cap\beta$ by the previous lemma. Therefore, $\mu([\sigma(P)]\boxtimes[\beta])=[P]\cap[\beta]$, and so the diagram commutes. ∎
3.3. K-homological maps $i^{!}$ and $i_{\ast}$ and $K$-theoretical map $i_{!}$
Let $i:\mathcal{Z}\to\mathcal{Y}$ be a closed subset (not necessarily an orbifold) of $\mathcal{Y}$ whose boundary is contained in that of $\mathcal{Y}$.
Recall that
$i_{\ast}^{fake}(H,\phi,T)=(H,\phi\circ i^{\ast},T).$
Conversely, we define
$i^{fake,!}(H,\phi,T)=(H,\phi\circ\tilde{i}_{!},T),$
where $\tilde{i}_{!}:\mathbb{C}(\mathcal{Z})\to\mathbb{C}(\mathcal{Y})$ is continuous and linear and is an extension for each function in $\mathbb{C}(\mathcal{Z})$.
Lemma 3.5.
Such an extension $\tilde{i}_{!}$ exists.
Proof.
One construction is as follows. We extend the real and imaginary parts separately in the same way so that the result is complex linear. We only need to extend functions on $\mathcal{Z}$ with maximum absolute value $1$. We modify the proof of Tietze’s extension theorem by incorporating a distance $d$ on $\mathcal{Y}$. For each $f$, in the first step, we set $f_{0}=f$,
$\begin{array}[]{ll}\mathcal{Z}_{k,+}&=\overline{\{x\in\mathcal{Z}:f_{k}(x)>(\frac{1}{3})^{k}\}},\\[2.84526pt] \mathcal{Z}_{k,-}&=\overline{\{x\in\mathcal{Z}:f_{k}(x)<-(\frac{1}{3})^{k}\}},\end{array}$
and define
$g_{k}(x)=\frac{-d(x,\mathcal{Z}_{k,+})+d(x,\mathcal{Z}_{k,-})}{d(x,\mathcal{Z}_{k,+})+d(x,\mathcal{Z}_{k,-})}\Bigl{(}\frac{1}{3}\Bigr{)}^{k};$
$f_{k+1}=f_{k}-g_{k}.$
The extension is given by $\sum_{k>0}g_{k}$. ∎
Lemma 3.6.
$i^{fake,!}$ is independent of the extension $\tilde{i}_{!}$.
Proof.
Let $(H,\phi_{1,2},T)$ be the representatives we constructed from different extensions. Then
$\tilde{T}_{t}=\left[\begin{matrix}T\cos(\frac{\pi}{2}t)&\sin(\frac{\pi}{2}t)I\\[2.84526pt] \sin(\frac{\pi}{2}t)I&-T\cos(\frac{\pi}{2}t)\end{matrix}\right]$
defines a homotopy from $(H,\phi_{1},T)\ominus(H,\phi_{2},T)$ to a degenerate element. ∎
By definition, we have the following lemma.
Lemma 3.7.
The following diagram commutes for any inclusion $i:\mathcal{Z}\to\mathcal{Y}$.
[…]
From now on, assume that $\mathcal{Z}$ is a suborbifold. We recall the wrong‐way $K$-cohomological push-forward
defined in [4]. Let $\mathcal{N}_{\mathbb{C}}=B(\pi_{\mathcal{Z}}^{\ast}\mathcal{N}\otimes\mathbb{C})$, which is identified with $B^{\ast}\mathcal{N}\to B^{\ast}\mathcal{Z}$ by identifying the real part with the fibers of $\mathcal{N}\to\mathcal{Z}$ and the imaginary part with the corresponding cotangent vectors. We define the K-theoretical push-forward in the special case of $\mathcal{Y}=\mathcal{N}$ as
$i_{!}^{fake}:\alpha\to(\pi^{\ast}\alpha)\otimes(str_{1}\wedge^{\ast}\mathcal{N}_{\mathbb{C}}^{\ast}).$
The push-forward $i_{!}^{fake}$ in the general case is the composition of $i_{!}^{fake}$ for a tubular neighborhood and the map $K^{0,fake}(B^{\ast}\mathcal{N},\partial B^{\ast}\mathcal{N})\to K^{0,fake}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})$ given by extending a virtual bundle trivially outside $\mathcal{N}$. Moreover, $i_{!}^{fake}$ is independent of the tubular neighborhood chosen [3].
We define the genuine $i_{!}$, $i_{\ast}$, and $i^{!}$ sector‐wise. It is straightforward from the definition that
$(ij)_{\ast}=i_{\ast}j_{\ast},\ \ \ \ (ij)^{!}=j^{!}i^{!},\ \ \ \ (ij)_{!}=i_{!}j_{!}.$
3.4. Compatibility